OpenStudy (anonymous):

One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single? My answer is 576 Am I right?

OpenStudy (anonymous):

First, you need to find x, which is the total number of people who downloaded music that month. 1x = 0.6x + 384 1x - 0.6x = 384 0.4x = 384 0.4x/0.4 = 384/0.4 x = 960 Now you have to look back at the question, which says that 60% of people downloaded this song. Simply take 60% of the value of x. 960 x 0.6 = 576. There's your answer!

OpenStudy (anonymous):

Oh my gosh! Thank you so much! :D

OpenStudy (anonymous):

Welcome!